OpenStudy (aravindg):

a 4 kg block a is placed on top of a 8 kg block B which rests on a smooth table,A,just slips on B when a force of 12 newton is applied on a .what is maximum horizontal force f required to make the combination of a and b move together?if it is to be applied on B.

OpenStudy (anonymous):

F=mu(8+4)*9.81

OpenStudy (anonymous):

sooo?

OpenStudy (anonymous):

Fba=4u*9.81 F=12 u=3/9.8

OpenStudy (anonymous):

36 N for sure medal please

OpenStudy (anonymous):

jamal on the right track

OpenStudy (aravindg):

no u r wrong no medal